OpenStudy (anonymous):

A new softball dropped onto a hard surface from a height of 25 inches rebounds to about 2/5 the height on each successive bounce. (a) Write a function representing the rebound height for each bounce. (b) After how many bounces would a new softball rebound less than 1 inch? I tried to solve this on my own but I think it is wrong. Could someone please help? :)

OpenStudy (anonymous):

25, 25(2/5), 25(2/5)^2, 25(2/5)^3,... H(n) = 25(2/5)^n, where n is the number of bounce

OpenStudy (anonymous):

25(2/5)^n < 1

OpenStudy (anonymous):

A)f(x) = 25(0.4)x; 3 bounces. B)f(x) = 25(0.4)x; 4 bounces. C)f(x) = 0.4(25)x; 4 bounces. D)f(x) = 0.4(25)x; 3 bounces. I am not really sure, these are the answers.. I am sorry I am not very good at Math

OpenStudy (anonymous):

25(2/5)^4 = 0.64 < 1, so B

OpenStudy (anonymous):

Thank you Thank you Thank you!!! :D